OpenStudy (anonymous):

does H2SO4 dissociate completaly

OpenStudy (alphadxg):

Yes of course, H2SO4 >> 2H+ + SO42-

OpenStudy (vincent-lyon.fr):

Actually, it is rather subtle In water, the first ionisation of H2SO4 is complete : \(H_2SO_4+H_2O \rightarrow HSO_4^- +H_3O^+\) so \(H_2SO_4\) does not exist in water: in this sense, you can say it dissociates completely. BUT the hydrogensulfate ion only partially dissociates according to: \(HSO_4^-+H_2O \rightleftharpoons SO_4^{2-}+H_3O^+\) So \(H_2SO_4\) is never FULLY dissociated into \(SO_4^{2-}+2H_3O^+\)

OpenStudy (anonymous):

yes because of H2SO4 is a strong acid and strong acid completely ionize in water

OpenStudy (alphadxg):

@Vincent-Lyon.Fr Sulfuric Acid (H2SO4) is a Diprotic Acid. A Diprotic Acid can donate two Hydrogen Ions (protons) per molecule in an Aqueous Solution. Sulfuric Acid can donate one proton to form Hydrogen Sulfate (HSO4) or both to form Sulfate (SO4). "Di" means two ions can be released and "Protic" means the ions that are released are protons.

OpenStudy (aaronq):

@Vincent-Lyon.Fr is right. suppose you have 2 M \(H_2SO_4\) and you put it in pure water (pH=7). using the henderson hasselbalch equation: \(pH=pKa+log\dfrac{[A^-]}{[HA]}\) \(7=(-6.6)+log\dfrac{[A^-]}{[HA]}\) \(\rightarrow\) \(\dfrac{[A^-]}{[HA]}=39810717055349.7\) there is 99.99999999999749% \(HSO_4^{-} \)and \(2.51*10^{-12}\)% \(H_2SO_4\) so you can assume the first proton is 100% dissociated. the pH is now \(pH=-log(2)=-0.3\) \(-0.3=(1.99)+log\dfrac{[A^-]}{[HA]}\) \(\rightarrow\) \(\dfrac{[A^-]}{[HA]}=0.00513\) so there is 0.5% \(SO_4^{2-} \)and 95.5% \(HSO_4^-\) present when you use 2M sulfuric acid.

OpenStudy (alphadxg):

@aaronq @Vincent-Lyon.Fr Well, it does dissociate completely in the first ionization, but I guess your right, not in the second ionization.